OpenStudy (anonymous):

How is the graph of log (x – 5) translated from the graph of log x? shifted right 5 units shifted down 5 units shifted up 5 units shifted left 5 units

OpenStudy (anonymous):

\[f(x)=\log(x-a)\]is shifted a units from\[f(x)=\log(x)~~.\]

OpenStudy (anonymous):

is shifted a units right from f(x)=log(x)

OpenStudy (anonymous):

so the answer would be shifted right?